# The General Position Problem in Graph Theory: A Survey

### v3

@@ 5.3 Colouring problems
Graph colourings in which each colour class is required to have some special property are very common in graph theory. To give a very few examples, in proper colourings (the foundational colouring problem) each colour class must constitute an independent set, in cocolourings each colour class must induce either a clique or an independent set, for 222-distance colourings any pair of vertices with the same colour must be at distance at least two, and the domatic number of a graph is the largest possible number of colours when each colour class is a dominating set.
A colouring version of Dudeney’s geometrical puzzle was considered in the short note [124]. In this article Wood shows that if $p$ is the smallest prime $\geq n$, then the $n\times n$ chessboard can be covered by $n+p-1$ sets of pawns of different colours such that each colour class has the no-three-in-line property, from which it follows that for any $\epsilon>0$ and sufficiently large $n$ at most $(2+\epsilon)n$ colours suffice. As far as we are aware the multiplicative constant has not been improved on since 2004. A more general colouring conjecture for any collection of points in the plane is presented in [93] and some progress towards this is contained in the thesis [92].
However, it seems that the first person to suggest investigating graph colourings in which each colour class is a general position set was E. Sampathkumar, and this idea was taken up in the recent paper [19]. The smallest number of colours needed to colour $V(G)$ such that each colour class is in general position is the $\operatorname{gp}$-chromatic number of $G$ and is written $\chi_{\operatorname{gp}}(G)$. For an example, in Figure 4 the red set on the left and the blue set on the right together constitute a $\operatorname{gp}$-colouring of the Petersen graph. Some simple bounds on the $\operatorname{gp}$-chromatic number are the following.
@@
[19, Lemma 3.2] If $G$ is a graph, then
@@
[19, Lemma 3.6] If $G$ is a connected graph, then
@@
[19, Theorem 4.10] There exists a diamond-free graph $G$ with $\chi_{\operatorname{gp}}(G)=a$ and $\gamma_{\rm t}(G)=b$ if and only if $1\leq a\leq b$.
@@
[19, Theorem 5.1] If $r\geq 1$ and $n_{1}\geq\dots\geq n_{r}$, then
@@
[19, Theorem 5.3] If $n\geq 5$, then $\chi_{\operatorname{gp}}(K(n,2))=n-3$.
@@
[19, Theorem 4.10] If $n\geq 3$, then
$\chi_{\operatorname{gp}}(L(K_{n}))=\begin{cases}\frac{n}{2}+1;&n\in\{6,12\},\\ \frac{n+1}{2};&n\equiv 1,5\bmod 6,\\ \frac{n}{2};&n\equiv 2,4\bmod 6,\text{ or }n\equiv 0\bmod 6\text{ and }n\geq 18,\\ \frac{n-1}{2};&n\equiv 3\bmod 6.\\ \end{cases}$
@@
[19, Theorem 7.4] If $n\geq 4$, then $\chi_{\operatorname{gp}}(P_{4}\operatorname{\,\square\,}P_{n})=n+1$.
@@
Figure 7: A $\operatorname{gp}$-colouring of a 4×64$\times$ 64 × 6 Cartesian grid
@@
[19, Theorem 7.2] If $n\geq 3$, then
$\chi_{\operatorname{gp}}(P_{2}\operatorname{\,\square\,}P_{n})=\begin{cases}2r;&n=3r,\\ 2r+1;&n=3r+1,\\ 2r+2;&n=3r+2.\\ \end{cases}$
@@
[…]
@@
[19, Theorem 4.12] For $a\geq 2$, the largest size of a graph with $\chi_{\operatorname{gp}}(G)=a$ is at least (${n\choose 2}-\frac{(a-1)a(a+1)}{6}$ for $n\geq\frac{a(a+1)}{2}$ and the largest size of a graph with $\chi_{\operatorname{mp}}(G)=a$ is at least (${n\choose 2}-(a-2)(2a-1)$ for $n\geq 2a$.
@@
[19, Conjecture 9.3] If $G$ is a graph and $\pi\in\{\operatorname{gp},\operatorname{mp}\}$, then $\chi_{\pi_{\rm i}}(G)+\chi_{\pi_{\rm i}}(\overline{G})\leq n+1$.
@@ 5.4 Fractional position problems
[44, Theorem 1.1] If $G$ is a graph, then $\operatorname{gp}_{f}(G)\geq\frac{2n(G)}{\operatorname{diam}(G)+1}$.
@@
[44, Theorem 1.2] If $n\geq 2$ then $\operatorname{gp}_{f}(K_{n})=n$, and if $n\geq 3$, then
@@
[44, Theorem 1.3] If $T$ is a tree, then $\operatorname{gp}_{f}(T)=\operatorname{gp}(T)=\ell(T)$.
@@
[44, Theorem 1.4] If $n\geq 1$, then $\operatorname{gp}_{f}(Q_{n})=\frac{2^{n+1}}{n+1}$.
[…]
@@
Lower 4 $\frac{24}{5}$ 6 $\frac{64}{9}$ $\frac{42}{5}$ $\frac{104}{11}$ $\frac{32}{3}$ $\frac{176}{15}$ 131313 $\frac{240}{17}$ $\frac{46}{3}$ $\frac{312}{19}$ $\frac{194}{11}$ $\frac{432}{23}$ 20
@@ 5.5 Games on graphs
[65, Theorem 2.2] Let $G$ be a graph. Then the following holds.
@@
[20, Theorem 2.4] Let $G$ be a graph. Then the following holds.
@@
[65, Theorem 2.3] If $k\geq 2$ and $n_{i}\geq 2$, $i\in[k]$, then A wins the $\operatorname{gp}$-achievement game on $K_{n_{1},\ldots,n_{k}}$ if and only if $k$ is odd and at least one $n_{i}$ is odd.
@@
[65, Theorem 2.5] Let $G$ be a bipartite graph. Then A wins the $\operatorname{gp}$-achievement game on $G$ if and only if the number of isolated vertices in $G$ is odd.
As an application of Theorem 5.40, the $\operatorname{gp}$-avoidance game is solved in [20, Theorem 2.5] for the complete multipartite graphs.
@@
[20, Theorem 2.6] If $n\geq 1$ and $m\geq 3$, then B wins the $\operatorname{gp}$-avoidance game on $W_{n,m}$ if and only if $m\geq 4$.
@@
[65, Theorem 3.6] Let $G$ be a connected graph and let $H$ be a connected bipartite graph with at least one edge. Then B wins the $\operatorname{gp}$-achievement game on $G\operatorname{\,\square\,}H$.
@@
[65, Theorem 3.7] If $n,m\geq 2$, then A wins the $\operatorname{gp}$-achievement game on $K_{n}\operatorname{\,\square\,}K_{m}$ if and only if both $n$ and $m$ are odd.
As remarked in [65, Corollary 3.5], if $n$ is even and $G$ is a connected graph, then B wins the $\operatorname{gp}$-achievement game on $K_{n}\operatorname{\,\square\,}G$. On the other hand, if $n$ is odd and $G$ is a connected graph, the outcome of the $\operatorname{gp}$-achievement game on $K_{n}\operatorname{\,\square\,}G$ appears to be difficult. This statement is justified by the proof of the following result.
@@
[65, Theorem 3.8] If $m\geq 3$, then A wins the $\operatorname{gp}$-achievement game on $K_{3}\operatorname{\,\square\,}C_{m}$ if and only if $m\in\{3,5\}$.
@@
[65, Theorem 4.3] If $G$ is a connected graph, then B wins the $\operatorname{gp}$-achievement game on $G\circ K_{n}$ if and only if B wins on $G$ or $n$ is even.
@@
[20, Theorem 4.4] If $n,m\geq 2$, then B wins the $\operatorname{gp}$-avoidance game on $K_{n}\operatorname{\,\square\,}K_{m}$ if and only if either $n=2$ and $m$ is odd, or $n=3$ and $m$ is even.
@@
[20, Theorem 4.5] If $n\geq 3$ and $m\geq 2$, then B wins the $\operatorname{gp}$-avoidance game on $P_{n}\operatorname{\,\square\,}P_{m}$.
@@
[20, Theorem 4.7] If $n\geq 3$ and $m\geq 2$, then B wins the $\operatorname{gp}$-avoidance game on $C_{n}\operatorname{\,\square\,}P_{m}$ if and only if $n$ is odd.
@@
[20, Theorem 5.4] If $G$ is a connected graph and $n\geq 1$, then B wins the $\operatorname{gp}$-avoidance game on $G\circ K_{n}$ if and only if B wins the $\operatorname{gp}$-avoidance game on $G$ and $n$ is odd.
@@
[…]
@@ 5.5.1 Complexity of games
The algorithmic complexity of deciding the winning player of the general position achievement/avoidance games of a given graph was studied in [20]. It was proved that the complexity of deciding the winning player of the general position achievement/avoidance games is in the class of PSPACE-complete problems even in graphs with diameter at most 444. Formally, they consider the games as decision problems: Given a graph $G$, does player A have a winning strategy?
@@
[20, Theorem 3.2] The game $\operatorname{gp}$-achievement is PSPACE-complete even on graphs with diameter at most four.
@@
[20, Theorem 3.3] The game $\operatorname{gp}$-avoidance is PSPACE-complete even in graphs with diameter at most four.
@@
[70, Conjecture 5.5] The decision versions of the B- and B’-games are PSPACE-complete.
@@ 6 Open problems
[…]
  • [124, Problem 2] What is the smallest value of $c$ such that $(c+\epsilon)n$ colours suffice for any $\epsilon>0$ and sufficiently large $n$ to cover an $n\times n$ chessboard with coloured pawns, such that each colour class has the no-three-in-line property? It is known that $\frac{1}{2}\leq c\leq 2$.
@@
  • [71, Problem 4.8] Is it true that if $G$ and $H$ are arbitrary connected graphs, then $\operatorname{gp}(G\boxtimes H)=\operatorname{gp}(G)\operatorname{gp}(H)$?
@@
  • [19, Conjecture 9.3] Is it true that if $G$ is a graph and $\pi\in\{\operatorname{gp},\operatorname{mp}\}$, then $\chi_{\pi_{\rm i}}(G)+\chi_{\pi_{\rm i}}(\overline{G})\leq n(G)+1$?
  • Prove that the detour position number of a graph is bounded above by $n-D+1$, where $D$ is the detour diameter.
  • [77, Conjecture 3.3] Does every graph contain a terminal set?
  • [77, Conjecture 2.10] Is it true that $\operatorname{gp}^{-}(G\operatorname{\,\square\,}H)\geq\min\{\operatorname{gp}^{-}(G),\operatorname{gp}^{-}(H)\}$ for any graphs $G,H$?
@@
  • [122, Conjecture 4.5] Prove that for any $n\geq 11$ there is a circulant graph with order $n$, monophonic position number two and diameter two.
@@
  • [19, Problem 9.2] (Packing problem) For a connected graph $G$, what is the largest number of disjoint maximal general position sets contained in $G$?
@@
  • [70, Conjecture 11] What is the complexity of the decision version of the Builder/Blocker general position game?

### v4

@@ 5.3 Colouring problems
Graph colourings in which each colour class is required to have some special property are very common in graph theory. To give a very few examples, in proper colourings (the foundational colouring problem) each colour class must constitute an independent set, in cocolourings each colour class must induce either a clique or an independent set, for $2$-distance colourings any pair of vertices with the same colour must be at distance at least two, and the domatic number of a graph is the largest possible number of colours when each colour class is a dominating set.
In discrete geometry colourings with a no-three-in-line property has long been an active area of research (see [9] for some references). A colouring version of Dudeney’s geometrical puzzle was considered in the short note [142]. In this article Wood shows that if $p$ is the smallest prime $\geq n$, then the $n\times n$ chessboard can be covered by n+p1n+p-1 sets of pawns of different colours such that each colour class has the no-three-in-line property, from which it follows that for any $\epsilon>0$ and sufficiently large $n$ at most $(2+\epsilon)n$ colours suffice. 22 years later, this was improved to $(1+\epsilon)n$ colours for any $\epsilon>0$ and sufficiently large $n$ in [9] as a by-product of new results on no-three-in-line (or arc-proper) colourings in finite geometries. A more general colouring conjecture for any collection of points in the plane is presented in [104] and some progress towards this is contained in the thesis [103].
It seems that the first person to suggest investigating graph colourings in which each colour class is a general position set was E. Sampathkumar, and this idea was taken up in the recent paper [22]. The smallest number of colours needed to colour $V(G)$ such that each colour class is in general position is the $\operatorname{gp}$-chromatic number of $G$ and is written $\chi_{\operatorname{gp}}(G)$. For an example, in Figure 4 the red set on the left and the blue set on the right together constitute a $\operatorname{gp}$-colouring of the Petersen graph. Some simple bounds on the $\operatorname{gp}$-chromatic number are the following.
@@
[22, Lemma 3.2] If $G$ is a graph, then
@@
[22, Lemma 3.6] If $G$ is a connected graph, then
@@
[22, Theorem 4.10] There exists a diamond-free graph $G$ with $\chi_{\operatorname{gp}}(G)=a$ and $\gamma_{\rm t}(G)=b$ if and only if $1\leq a\leq b$.
@@
[22, Theorem 5.1] If $r\geq 1$ and $n_{1}\geq\dots\geq n_{r}$, then
@@
[22, Theorem 5.3] If $n\geq 5$, then $\chi_{\operatorname{gp}}(K(n,2))=n-3$.
@@
[22, Theorem 4.10] If $n\geq 3$, then
[…]
@@
[22, Theorem 7.4] If $n\geq 4$, then $\chi_{\operatorname{gp}}(P_{4}\operatorname{\,\square\,}P_{n})=n+1$.
@@
Figure 7: A $\operatorname{gp}$-colouring of a $4\times 6$ Cartesian grid
@@
[22, Theorem 7.2] If $n\geq 3$, then
[…]
@@
[22, Theorem 8.2] Deciding if a graph $G$ has a $\operatorname{gp}$-colouring with $\leq k$ colours is NP-complete even for instances $(G,k)$ with $\operatorname{diam}(G)=2$ and $k=3$.
@@
[22, Theorem 4.12] For $a\geq 2$, the largest size of a graph with $\chi_{\operatorname{gp}}(G)=a$ is at least ${n\choose 2}-\frac{(a-1)a(a+1)}{6}$ for $n\geq\frac{a(a+1)}{2}$ and the largest size of a graph with $\chi_{\operatorname{mp}}(G)=a$ is at least ${n\choose 2}-(a-2)(2a-1)$ for $n\geq 2a$.
@@
[22, Conjecture 9.3] If $G$ is a graph and $\pi\in\{\operatorname{gp},\operatorname{mp}\}$, then $\chi_{\pi_{\rm i}}(G)+\chi_{\pi_{\rm i}}(\overline{G})\leq n(G)+1$.
@@ 5.4 Fractional position problems
[49, Theorem 1.1] If $G$ is a graph, then $\operatorname{gp}_{f}(G)\geq\frac{2n(G)}{\operatorname{diam}(G)+1}$.
@@
[49, Theorem 1.2] If $n\geq 2$ then $\operatorname{gp}_{f}(K_{n})=n$, and if $n\geq 3$, then
@@
[49, Theorem 1.3] If $T$ is a tree, then $\operatorname{gp}_{f}(T)=\operatorname{gp}(T)=\ell(T)$.
@@
[49, Theorem 1.4] If $n\geq 1$, then $\operatorname{gp}_{f}(Q_{n})=\frac{2^{n+1}}{n+1}$.
The majority of the paper focusses on the square grid $P_{n}\operatorname{\,\square\,}P_{n}$. Finding the fractional general position number of this graph turns out to be a deep problem. The following lower bound, which is the best known construction for even $n$, is obtained by assigning label 0 to every vertex at distance <t<t from a corner and label $\frac{2}{2n-2t-1}$ to the remaining vertices, for some integer in the range $1\leq t\leq\frac{n}{2}$.
@@
Lower 4 $\frac{24}{5}$ 6 $\frac{64}{9}$ $\frac{42}{5}$ $\frac{104}{11}$ $\frac{32}{3}$ $\frac{176}{15}$ $13$ $\frac{240}{17}$ $\frac{46}{3}$ $\frac{312}{19}$ $\frac{194}{11}$ $\frac{432}{23}$ 20
@@
We note that [9] introduces a fractional version of arc-proper colourings in finite geometry. This suggests investigating fractional versions of the problems discussed in the preceding subsection.
@@ 5.5 Games on graphs
[75, Theorem 2.2] Let $G$ be a graph. Then the following holds.
@@
[23, Theorem 2.4] Let $G$ be a graph. Then the following holds.
@@
[75, Theorem 2.3] If $k\geq 2$ and $n_{i}\geq 2$, $i\in[k]$, then A wins the $\operatorname{gp}$-achievement game on $K_{n_{1},\ldots,n_{k}}$ if and only if $k$ is odd and at least one $n_{i}$ is odd.
@@
[75, Theorem 2.5] Let $G$ be a bipartite graph. Then A wins the $\operatorname{gp}$-achievement game on $G$ if and only if the number of isolated vertices in $G$ is odd.
As an application of Theorem 5.40, the $\operatorname{gp}$-avoidance game is solved in [23, Theorem 2.5] for the complete multipartite graphs.
@@
[23, Theorem 2.6] If $n\geq 1$ and $m\geq 3$, then B wins the $\operatorname{gp}$-avoidance game on $W_{n,m}$ if and only if $m\geq 4$.
@@
[75, Theorem 3.6] Let $G$ be a connected graph and let $H$ be a connected bipartite graph with at least one edge. Then B wins the $\operatorname{gp}$-achievement game on $G\operatorname{\,\square\,}H$.
@@
[75, Theorem 3.7] If $n,m\geq 2$, then A wins the $\operatorname{gp}$-achievement game on $K_{n}\operatorname{\,\square\,}K_{m}$ if and only if both $n$ and $m$ are odd.
As remarked in [75, Corollary 3.5], if $n$ is even and $G$ is a connected graph, then B wins the $\operatorname{gp}$-achievement game on $K_{n}\operatorname{\,\square\,}G$. On the other hand, if $n$ is odd and $G$ is a connected graph, the outcome of the $\operatorname{gp}$-achievement game on $K_{n}\operatorname{\,\square\,}G$ appears to be difficult. This statement is justified by the proof of the following result.
@@
[75, Theorem 3.8] If $m\geq 3$, then A wins the $\operatorname{gp}$-achievement game on $K_{3}\operatorname{\,\square\,}C_{m}$ if and only if $m\in\{3,5\}$.
@@
[75, Theorem 4.3] If $G$ is a connected graph, then B wins the $\operatorname{gp}$-achievement game on $G\circ K_{n}$ if and only if B wins on $G$ or $n$ is even.
@@
[23, Theorem 4.4] If $n,m\geq 2$, then B wins the $\operatorname{gp}$-avoidance game on $K_{n}\operatorname{\,\square\,}K_{m}$ if and only if either $n=2$ and $m$ is odd, or $n=3$ and $m$ is even.
@@
[23, Theorem 4.5] If $n\geq 3$ and $m\geq 2$, then B wins the $\operatorname{gp}$-avoidance game on $P_{n}\operatorname{\,\square\,}P_{m}$.
@@
[23, Theorem 4.7] If $n\geq 3$ and $m\geq 2$, then B wins the $\operatorname{gp}$-avoidance game on $C_{n}\operatorname{\,\square\,}P_{m}$ if and only if $n$ is odd.
@@
[23, Theorem 5.4] If $G$ is a connected graph and $n\geq 1$, then B wins the $\operatorname{gp}$-avoidance game on $G\circ K_{n}$ if and only if B wins the $\operatorname{gp}$-avoidance game on $G$ and $n$ is odd.
@@
An unexpectedly interesting question is: does it matter who goes first? Based on the domination game and its ‘Continuation Principle’, see [18, 19], intuition may suggest that the order of the players does not make a big difference. This turns out to be completely false. We call a pair $(a,b)$ realisable if there is a graph $G$ with $\operatorname{gp}_{\rm g}(G)=a$ and $\operatorname{gp}_{\rm g}^{\prime}(G)=b$. Proving that pairs with $2\leq a\leq b$ are realisable is quite simple, but for $a<b$ the problem is harder. A construction involving identifying 4-cycles and an odd number of 3-cycles along an edge shows that $(a,b)$ is realisable if $a>b$ and $b\geq 3$ is odd. Some such pairs with even $b$ were also found. The article therefore raised the question: is every pair $(a,b)$ with $a>b>2$ and even $b$ realisable? This question was completely settled in the affirmative in a forthcoming article [48]. Finding the smallest graphs that realise a given pair also seems a worthwhile problem.
@@ 5.5.1 Complexity of games
The algorithmic complexity of deciding the winning player of the general position achievement/avoidance games of a given graph was studied in [23]. It was proved that the complexity of deciding the winning player of the general position achievement/avoidance games is in the class of PSPACE-complete problems even in graphs with diameter at most $4$. Formally, they consider the games as decision problems: Given a graph $G$, does player A have a winning strategy?
@@
[23, Theorem 3.2] The game $\operatorname{gp}$-achievement is PSPACE-complete even on graphs with diameter at most four.
@@
[23, Theorem 3.3] The game $\operatorname{gp}$-avoidance is PSPACE-complete even in graphs with diameter at most four.
@@
[80, Conjecture 5.5] The decision versions of the B- and B’-games are PSPACE-complete.
@@ 6 Open problems
  • Find a value of $n$ for which $2n$ pawns cannot be placed in general position in the No-Three-In-Line Problem, or show that $2n$ pawns is always possible (perhaps for large n) (Hard).
  • [9, 142] What is the smallest value of $c$ such that $(c+\epsilon)n$ colours suffice for any $\epsilon>0$ and sufficiently large $n$ to cover an $n\times n$ chessboard with coloured pawns, such that each colour class has the no-three-in-line property? It is now known that $\frac{1}{2}\leq c\leq 1$.
@@
  • [81, Problem 4.8] Is it true that if $G$ and $H$ are arbitrary connected graphs, then $\operatorname{gp}(G\boxtimes H)=\operatorname{gp}(G)\operatorname{gp}(H)$?
@@
  • [22, Conjecture 9.3] Is it true that if $G$ is a graph and $\pi\in\{\operatorname{gp},\operatorname{mp}\}$, then $\chi_{\pi_{\rm i}}(G)+\chi_{\pi_{\rm i}}(\overline{G})\leq n(G)+1$?
  • Prove that the detour position number of a graph is bounded above by nD+1n-D+1, where $D$ is the detour diameter.
  • [88, Conjecture 3.3] Does every graph contain a terminal set?
  • [88, Conjecture 2.10] Is it true that $\operatorname{gp}^{-}(G\operatorname{\,\square\,}H)\geq\min\{\operatorname{gp}^{-}(G),\operatorname{gp}^{-}(H)\}$ for any graphs $G,H$?
@@
  • [140, Conjecture 4.5] Prove that for any $n\geq 11$ there is a circulant graph with order $n$, monophonic position number two and diameter two.
@@
  • [22, Problem 9.2] (Packing problem) For a connected graph $G$, what is the largest number of disjoint maximal general position sets contained in $G$?
@@
  • [80, Conjecture 11] What is the complexity of the decision version of the Builder/Blocker general position game?